### v1

# Navigating Rifts in Human-LLM Grounding: Study and Benchmark

Abstract
Language models excel at following instructions but often struggle with the collaborative aspects of conversation that humans naturally employ. This limitation in grounding—the process by which conversation participants establish mutual understanding—can lead to outcomes ranging from frustrated users to serious consequences in high-stakes scenarios. To systematically study grounding challenges in human-LLM interactions, we analyze logs from three human-assistant datasets: WildChat, MultiWOZ, and Bing Chat. We develop a taxonomy of grounding acts and build models to annotate and forecast grounding behavior. Our findings reveal significant differences in human-human and human-LLM grounding: LLMs were three times less likely to initiate clarification and sixteen times less likely to provide follow-up requests than humans. Additionally, early grounding failures predicted later interaction breakdowns. Building on these insights, we introduce Rifts: a benchmark derived from publicly available LLM interaction data containing situations where LLMs fail to initiate grounding. We note that current frontier models perform poorly on Rifts, highlighting the need to reconsider how we train and prompt LLMs for human interaction. To this end, we develop a preliminary intervention that mitigates grounding failures.
Navigating Rifts in Human-LLM Grounding:
Study and Benchmark
Omar Shaikh†††thanks: Research performed during an internship at Microsoft., Hussein Mozannar⋄, Gagan Bansal⋄, Adam Fourney⋄, Eric Horvitz⋄ †Stanford University, ⋄Microsoft Research oshaikh@stanford.edu
1 Introduction
Language models used for conversational interaction are trained primarily to follow instructions Ouyang et al. (2022). But effective dialogue requires more than just instruction-following. Participants in conversation work together in a collaborative process, resolving ambiguities as they exchange ideas and achieve shared objectives. They cultivate common ground111We use grounding to refer to Clark’s formulation of the language, gestures, and signaling that participants in a conversation employ to establish and maintain an effective dialog with shared understanding among actors Clark (1996)—in distinction to “grounding” in multimodal grounding. through grounding processes: communicative behaviors aimed at establishing and confirming mutual comprehension Clark (1996). Grounding mechanisms guide how individuals anticipate and react to a conversation partner’s contributions (Clark and Schaefer, 1989). Speakers implement grounding through dialogue acts. They work to confirm or clarify assumptions, and ask follow-up questions. When grounding breaks down, participants employ repair strategies to resolve potential miscommunications.
In contrast, LLMs employed in conversational systems generate task-centric content with minimal grounding actions Shaikh et al. (2024). Failing to ground with users can be costly, with outcomes ranging from the common situation of frustrated users to that of serious consequences in high-stakes situations (Figure 1). The process by which humans build common ground offers a powerful metaphor for understanding human-computer interaction Brennan (2014). Ideal human-LLM grounding should allow for both human and machine initiatives, aimed at detecting misunderstandings and achieving mutual understanding Horvitz (1999).
We start by assessing the current state of grounding in human-LLM interactions, analyzing grounding behavior through real-world interaction logs. To examine human-LLM grounding, we first characterize interactions through a set of validated dialogue acts (§3). In synthesizing these grounding acts, we build on prior work in conversational analysis and dialog systems. We cover acts that communicate progress (acknowledgment, follow-up, etc.) and difficulty in grounding (repair, clarification, etc.), initiated by either a human or LLM.
Using a set of grounding acts, we construct models of human-LLM grounding, which we use to annotate and forecast grounding in conversation (§4-5). We construct and validate an LLM-based annotator, which enables efficient annotation of logs from publicly available interactions with ChatGPT (WildChat), data from a widely used commercial LLM service (Bing Chat), and Wizard-of-Oz-ed interactions with a human roleplaying as an AI (MultiWOZ). We additionally develop a grounding forecaster that predicts the presence of grounding acts in future turns. The forecaster enables a dialog intervention that can proactively prevent grounding difficulties in forthcoming turns.
Our analysis of human-LLM interaction data reveals significant asymmetries in initiating grounding: people are three times more likely to clarify and sixteen times more likely to issue follow-up requests compared to LLMs. In addition, we find that grounding failures occurring early in dialog cascade into higher likelihoods of downstream failure. Overall, there is substantial room for improvement in human-LLM interaction via addressing deficits in conversational grounding.
To systematically measure human-LLM grounding and test interventions, we introduce a new benchmark (§6). Rifts222Rifts can be accessed at this link: https://github.com/microsoft/rifts. is a curated set of 1.8K tasks—directly sourced from in-the-wild interaction logs—that require selective use of clarification and follow-up requests for interactive grounding. Most frontier models struggle with Rifts. While we propose an effective intervention with our forecaster, progress on Rifts will require rethinking how LLMs are trained to interact with people.
2 Related Work
Ambiguity and Common Ground
Disambiguating questions like “Did you see (the man (with the telescope))?” requires establishing common ground. While NLP benchmarks address ambiguity Min et al. (2020); Tamkin et al. (2022), they focus on well-defined cases (e.g., reference ambiguity). However, people engage LLMs for open-ended tasks—creative writing, cover letters—where correct answers aren’t predefined, making common ground crucial. Using naturalistic human-LLM interactions, we identify challenges in building common ground. To operationalize it Clark and Schaefer (1989); Clark (1996); Stalnaker (2002), we synthesize dialogue acts based on subdialogues Litman (1985); Litman and Allen (1987) and conversation structure Jefferson (1972).
Grounding in Dialog Systems
NLP systems from ELIZA onward have been designed to initiate some form of conversational grounding Weizenbaum (1966); Purver (2004); Li et al. (2023); Paranjape and Manning (2021). Decision-theoretic models have helped systems manage uncertainty about user goals Horvitz and Paek (2000a, b); Paek and Horvitz (2003), while multimodal approaches consider language and visual cues Pejsa et al. (2014). Human-LLM grounding is crucial for tasks including goal-coordination Bara et al. (2021); Mohanty et al. (2023); Fried et al. (2022); Li and Boyer (2015), planning Chu-Carroll and Carberry (1998); Lochbaum (1998), games Madureira and Schlangen (2023b); Shaikh et al. (2023), data retrieval Lu et al. (2023), improvisation Cho and May (2020), and design Vaithilingam et al. (2024). AI-initiated grounding improves conversation quality Zhou et al. (2022), enables human-AI collaboration Lin et al. (2023), and encorages humans to reflect on LLM outputs Park and Kulkarni (2023). Our work extends this by introducing methods and a benchmark for studying grounding in real-world human-LLM dialogue.
Proactive Mitigation of Grounding Failures
Research has explored using machine-learned models to predict and mitigate grounding failures in spoken dialog systems. By forecasting potential failures, models can guide proactive interventions. For example, in a call-routing system, predictions of downstream grounding issues were used to trigger early transfers to human operators, reducing user frustration and disengagement, like pressing keys to access human assistants or abandoning calls Horvitz and Paek (2007).
| Type | Act | Example |
| Instruction |
| Next Turn |
| Continuing the conversation with the next relevant turn. | ||
| Note: all other successful acts are a subset of Continue. | ||
| Acknowledge |
| Verbalizing understanding: “I see, O.K.” / repeating instruction. | ||
| Follow-up |
| Overresponse |
| Clarify |
| Repair |
| Reformulate |
| Restart |
LLMs and conversational grounding
Current LLMs appear to guess user intent and progress with assumptions of correct inferences rather than resolving uncertainties through grounding.333https://openai.com/blog/chatgpt This manifests as generations of over-informative responses Tsvilodub et al. (2023), refusal to handle ambiguity Abercrombie et al. (2023); Min et al. (2020); Gao et al. (2021), and overconfidence Mielke et al. (2022). Prior work demonstrated that LLM-powered conversational agents fail to generate appropriate grounding acts Shaikh et al. (2024); Lu et al. (2024). Rather than measuring human-LLM interaction through end-to-end evaluation Lee et al. (2022); Chiang et al. (2024), we consider discrete grounding acts in dialogues. Like Schneider et al. (2024) and Shaikh et al. (2024), we use prompted LLMs to classify these acts. While prior work has explored generating clarification requests through prompting Kuhn et al. (2022); Chen et al. (2023) and finetuning Andukuri et al. (2024); Zhang and Choi (2023); Hong et al. (2023); Gan et al. (2024), we examine grounding acts more broadly.
3 Human-LLM Grounding Acts
To measure grounding between people and LLMs, we curate a set of dialogue acts that serve as proxies for grounding. We outline our selected acts and discuss prior work motivating each act. Our typology builds on prior work in conversational grounding: Clark and Schaefer (1989) and Traum and Hinkelman (1992) outline a hierarchy of actions—including discourse acts—that are used by humans to ground with one another. Recent work has revisited conversational grounding in the context of LLMs Shaikh et al. (2024); Schneider et al. (2024), focusing on a subset of acts generated mainly by people (e.g, following-up, acknowledging understanding, and clarifying).
In contrast, we consider acts generated by both LLMs and people. Our selected acts serve as signals for effective grounding; we segment acts across communicated grounding outcomes. We focus on the construction of common ground, a potential grounding failure, or acknowledging . Using our typology, we can measure grounding outcomes with observable dialog acts during human-LLM interaction (illustrative examples in Table 1).
3.1
Addressing acts explicitly signal a potential misunderstanding. Here, participants are addressing this failure in an interaction.
occur when a participant repeats or restates their query because of a failure to ground. An utterance is a reformulation if the succeeding utterance from the same participant is semantically equivalent to the original. Reformulations are prevalent in search engine and information retrieval domains Lau and Horvitz (1999).
occur when users reset a conversation to improve understanding and achieve a successful dialog. They often follow significant misunderstandings, whether in the initial response or across multiple exchanges. Users may restart due to LLMs misinterpreting intent, ambiguity, sensitivity, or irrelevant context Shi et al. (2023), akin to search query retries after irrelevant results Lau and Horvitz (1999). Research on restarts includes user decisions to suspend a dialog and seek alternate solutions, such as transferring from AI-based systems to human operators before frustration escalates Horvitz and Paek (2007). Restarts are classified when an initial instruction is repaired or reformulated within 30 minutes Radlinski and Joachims (2005); Downey et al. (2007).
3.2
Grounding may also be uncertain between participants. Rather than clearly indicating success or failure, ambiguous acts represent strategies that participants use to—potentially inefficiently—prevent potential misunderstandings.
occur when a participant disambiguates an utterance from another participant; or when a participant proactively “clears up” misunderstandings. Clarifications often occur when a task is perceived as ambiguous (e.g. What did you mean by that?). Ginzburg and Cooper (2001); Purver et al. (2003b, a); Healey et al. (2011, 2003); Purver (2004); Stoyanchev et al. (2013); Madureira and Schlangen (2023a); Rahmani et al. (2023).
include more than what another participant reasonably asked for. Unlike Next Turn, which provides only expected information, overresponses also anticipate potential follow-ups. They often appear as overly verbose LLM-generated answers, a behavior contemporary reward models are criticized for encouraging by favoring longer responses Singhal et al. (2023).
3.3
Finally, advancing acts communicate that a participant understood past utterances from another participant in an interaction.
refer to the next conversational move made by a listener that is expected, given the prior turn(s) in a dialogue. Examples of relevant next turns include directly answering a question, expressing an opinion (agreeing or disagreeing), or apologizing. If no misunderstanding has occurred, a listener has moved on to the next relevant turn by default. Levinson et al. (1983); Sacks et al. (1978); Schiffrin (1987). Note that advancing grounding will initiate the Next Relevant Turn. We focus on two: follow-ups & acknowledgments.
elaborate on a prior utterance in an interaction. Unlike clarifications—which disambiguate or clarify—follow-ups seek additional information. Because follow-ups build on a prior utterance, they implicitly signal understanding of past utterances. Davis (1982); Graesser et al. (1995); Traum and Hinkelman (1992); Bunt et al. (2017).
explicitly signal understanding. These requests manifest either through explicit dialogue (e.g. I see; I understand; O.K.) or by repeating portions of another participant’s utterance (e.g. I can help you [write a story].) Unlike Reformulation, where a listener repeats to address failure, acknowledgment occurs when a speaker repeats to demonstrate understanding. Schegloff (1982); Sacks et al. (1978); Schiffrin (1987); Clark and Schaefer (1989); Cho and May (2020)
4 Data
To analyze collaborative grounding with LLM-based assistants, we draw from three English datasets consisting of dialogs between a human and an assistant. WildChat is a real-world Human-AI interaction dataset with interaction between people and several OpenAI models Zhao et al. (2024). User data was collected with consent, in exchange for free access to the models. We use the non-toxic version of WildChat, filter conversations to be in English, and sample one conversation from each user. Bing Chat, similar to WildChat, was collected from a large chat-based service used by millions of users, powered by OpenAI LLMs. Finally, MultiWOZ is a crowdsourced dataset of dialogue-based interaction with an assistant Budzianowski et al. (2018); Ramadan et al. (2018); Eric et al. (2019); Zang et al. (2020). In contrast to WildChat and Bing Chat, MultiWOZ contains human-human dialogs, with one human playing a “wizard-of-oz” role as the assistant. We use MultiWOZ 2.2 Zang et al. (2020), examining collaborative grounding acts on the validation and test splits for a subset of the tasks. While we observe similarities in terms of the tasks posed by humans across the three datasets, differences do exist, which makes direct comparisons difficult. In Appendix A.1, we outline the number of dialogs and messages in each dataset.
5 Modeling Human-LLM Grounding
Given a defined set of grounding acts and data drawn from logs of human-assistant interaction, we can build grounding models. We first build prompted classifiers that identify acts post-hoc and describe the status quo of human-LLM grounding (§5.1). Then, we train grounding forecasters, enabling forecasting of grounding acts given just the initial instruction (§5.2). With forecasters, we can identify tasks where grounding is critical and intervene when appropriate.
5.1 Classifying Human-LLM Grounding Acts
Method
We employ GPT-4o-mini to annotate grounding acts across a subset of our datasets. On subsets of our data, we observed nearly identical results using GPT-4o compared to GPT-4o-mini. Thus, we employ GPT-4o-mini for its affordability to allow for efficiency and reproducibility. Following Shaikh et al. (2024), we first encode our typology in a prompt (Appendix E.1) and label each turn in the conversations. To validate the accuracy of the approach, three authors annotated 10 dialogues (total of 108 messages) from each dataset and reached an average Cohen Kappa agreement of 0.71 across the datasets. Ties were broken through discussion, converging on a final selection of the majority label. Using this as a withheld test set, we find that Macro F-1 scores are reasonably high across datasets (0.75). A full table of results across labels is also in Appendix E.1.
Results
We observe significant differences between people and LLMs when initiating actions aimed at grounding. In datasets where LLMs serve as assistants, we observe that grounding acts are taken primarily by people. People repair and reformulate instructions at high rates; averaged across human-AI interaction datasets, and of human turns are labeled as reformulate and repair respectively (Figure 2). In contrast to human-LLM interactions, human-human interaction data (MultiWOZ) has fewer reformulate ( ) and repair ( ) acts initiated by humans.
Session restarts serve as a final fallback when repair or reformulation fail to address a communicated failure. We focus on users with multiple interactions on WildChat (the only dataset that includes user IDs) and identify if a session begins with a repair or reformulation of an earlier instruction issued within the last 30 minutes. of sessions are restarts of a session in the last 30 minutes—exceeding the rate of repair in a conversation.
Before a human ends up addressing grounding, an ideal LLM assistant would have proactively engaged in clarification. However, we find that LLMs clarify at significantly lower rates (, t-test) compared to humans repairing. In fact, the opposite occurs: people clarify LLM outputs ( ) 3 times as much as LLMs clarify user instructions ( ; Fig. 3). In contrast, humans and wizard-of-oz-ed assistants clarify at similar rates (MultiWOZ; human user v.s. human assistant).
Beyond repairing/reformulating, people regularly ask directed follow-up questions when signaling at successful grounding (Figure 4). Across WildChat and Bing Chat, users ask times more follow-ups than LLM assistants. This disparity is less apparent in human-human interaction data, where human users only follow up times more. Instead of generating follow-ups, LLM assistants regularly over-respond ( of assistant turns), generating verbose responses that answer more than what the user asked for. Humans rarely overrespond—both when interacting with LLMs ( ) or when roleplaying an assistant ( ).
Early grounding patterns also have rippling effects: we find evidence of compounding advancing and addressing patterns as a conversation progresses (Fig. 5). Let denote the message at turn , which can be an advancing grounding act or addressing where and denote the sets of advancing and addressing grounding acts. The likelihood of an addressing utterance grounding appearing in the first turn on WildChat, and on Bing Chat. We observe a compounding of addressing grounding in human-assistant conversations. If we assume that the turn before was addressing, we find that the likelihood of another failed interaction triples: for Wildchat and for Bing Chat. A similar effect appears for advancing grounding: early signs of effective conversational grounding similarly snowballs, with on Wildchat and on Bing Chat.
5.2 Forecasting a User’s Grounding Acts
The GPT-based annotator (from §5.1) annotates conversations with grounding acts in a post-hoc fashion; we only identify opportunities for grounding after they happen. In contrast, curating tasks where grounding failures emerge requires predicting the likelihood of future grounding patterns. We focus on WildChat as it is the only publicly availible human-LLM interaction dataset in our analysis.
| Category | Prompt |
| Advancing | Write a Main heading about a brand name FFF Digital , which is a digital marketing agency |
| Suggest a name for a technical blog consisting of five characters at most, which is compatible with SEO | |
| 1 week out from my powerlifting meet and i’m not prepared […] what should i do? [omitted context] | |
| Addressing | Blackburn rovers vs West Bromwich albion prediction |
| I need to remove a heart | |
| [snippets of code with no prompt] | |
| Ambiguous | What causes tailbone pain? |
| My friend not want to help me, what to [do] with him? | |
| What happens when someone quits a job without having another one lined up? | |
| No Grounding | convert rust String to clap::builder::Str |
| Generate a full harvard references section for the following report: [REPORT] | |
| Join now, Supplier! or Supplier, Join us! which one is better? |
Method
At turn , our goal is to forecast grounding act associated with user message . We achieve this by repurposing conditional training—concretely, we append each user message with a forecasting grounding token if the future turn contains a grounding act. We use our validated grounding acts labeler (§5.1) to obtain , using the high-level grounding categories as labels (e.g. , , and ).
We finetune Llama-3.1-8B on sequences of form where is the assistant response. At inference time, we provide and analyze the likelihoods (i.e. logits) of our tokens . We detail hyperparameters used in the training process in Appendix B. Beyond our finetuned forecaster, we also evaluate as a baseline few-shot prompted GPT-4o-mini (details in Appendix E.2).
Results
We find that our GPT-4o-mini few-shot baseline performs near-random, with Macro ROC AUC = 0.51. This result is not very surprising as forecasting is a challenging task: we must predict grounding acts without directly observing assistant responses . Our finetuned forecaster, however, performs significantly better (0.61). Full experimental results are in Appendix B. In the next section, we draw representative samples from the forecaster, constructing a benchmark where users are (un)likely to initiate a grounding process.
6 Rifts: A Grounding Benchmark
We showed that LLMs fail to generate grounding acts in two settings. They rarely:
Clarify goals to reduce the rate at which a user is likely to address grounding.
Follow-up to advance grounding, instead of relying on the user to take initiative.
To characterize these behaviors across multiple LLMs and evaluate interventions, we introduce a new benchmark, Rifts. Rifts consists of tasks designed to test if LLMs can generate grounding acts when needed and withhold appropriately.
Dataset Details
Rifts consists of a final combined set of 1740 tasks (split counts in Appdx. Table 8). Each task in Rifts is a prompt drawn from an initial instruction in WildChat. Tasks are stratified based on how the user is predicted to continue the conversation: namely, are users predicted to grounding, a failure, or clarify an following an LLMs response (examples in Table 2). For these tasks, we would expect an LLM to take initiative—clarifying or following up appropriately. In addition, we include tasks where the user is expected to do none of the above, perhaps by switching the topic or ending the interaction altogether. Here, no grounding is required.
Curation Process
We construct Rifts by filtering prompts from WildChat, using the predicted grounding act of the forecaster from §5.2. Forecaster predictions inform us on whether a clarification or followup action will be required in the conversation. We first filtered correctly predicted tasks from forecasters trained444Why train on each split? To ensure fair evaluation, we create separate forecasters for train / val / test. This prevents leakage and enables researchers to develop interventions using the train forecaster before evaluating with the test forecaster. on each split (train / val / test). For each grounding category, we then extracted the top 150 tasks with the highest likelihood of generating advancing, addressing, or ambiguous forecasting tokens. In other words, we repurpose our forecaster to curate representative tasks across each grounding act, sorting by the logit associated with each forecasting token. In addition, we sample 150 tasks that have a high likelihood of generating no forecasting token, capturing tasks that do not require initiative. Finally, we apply basic quality controls (see Appendix C).
Evaluating LLMs
Rifts simplifies evaluation for any assistant model . Consider the two failure modes where LLMs do not take initiative: clarify and followup. Given a task from Rifts in the category, we would prefer to proactively generate a followup. On the other hand, for tasks in or , we would expect to generate clarification questions. In instances where we forecast no act from the user, we do not wish to see the model inefficiently engage in grounding activity.
To evaluate performance, we take an initial instruction from Rifts, and sample the next turn from . We then label with our validated grounding acts annotator (§5.1). To benchmark , we evaluate if the generated response clarifies/follows-up when appropriate. Concretely, we instantiate our two failure modes (clarify, followup) in the following EVAL() function and report an overall accuracy score:
| Model | Variant | Rifts Accuracy |
| GPT | 4o | 25.26 3.54 |
| 4o-mini | 24.48 3.51 | |
| o3-mini | 25.26 3.54 | |
| Claude | Sonnet 3.5 | 26.95 3.57 |
| Opus 3 | 24.57 3.51 | |
| Llama 3.1 | 8B Instruct | 24.22 3.49 |
| 70B Instruct | 23.88 3.47 | |
| Llama 3.1 | 8B + GROUND | 54.48 2.45 |
Off-the-shelf models struggle
We evaluate a handful of open- and closed-source models on Rifts’ test set: OpenAI’s GPT-4(o) series, Anthropic’s Claude Sonnet 3.5 / Opus 3, and Llama 3.1 8 / 70B (Table 3). We find that all off-the-shelf instruction-following models (avg. 23.23% acc.) perform worse than random (33%). All of our evaluated LLMs perform near perfectly for tasks that require no grounding initiative (No Grounding category, 96.09%); this is unsurprising given instruction-following. However, LLMs fail to take appropriate initiative for any of the remaining categories (2.22% of the time, Table 7). Reasoning-tuned models don’t help either: o3-mini regularly begins reasoning without verifying grounding.
A simple intervention
To improve grounding capabilities, we turn again to our forecasters (§5.2). Depending on the train forecaster’s prediction, we can selectively add a prompt (+ GROUND) that instructs the LLM to ask follow-up questions or request clarification (prompts in Appendix E.3). Concretely, we append a clarification prompt if our forecaster predicts or ; or a follow-up prompt if our forecaster predicts . With this intervention, Llama 3.1 8B outperforms all other models by at least 32%. Still, our intervention is far from perfect. Rifts opens avenues for benchmarking new interventions, enabling easy evaluation of grounding capabilities in future work.
7 Discussion and Conclusion
We characterized (§3) and measured (§4-5) inadequate grounding in human-LLM interaction; and proposed a benchmark (§6) to assess this gap. Several directions emerge:
Towards LLMs that Initiate Grounding
Decision-theoretic methods could guide when and how LLMs initiate grounding actions, based on inferred uncertainties in mutual understanding (see Horvitz (1999); Mozannar et al. (2024)). Instruction tuning could be revised to incorporate grounding, and our forecaster could serve as a reward model in RLHF Ouyang et al. (2022). System prompts and dialogue management show promise, including prompts to disambiguate user intentions Chen et al. (2023).
Benchmarking Human-LLM Grounding
Building models that ground effectively with humans across a range of tasks requires effective benchmarks. Rifts supports comparative analyses, enabling discussion on grounding competencies of new LLMs and interventions.
Limitations
We considered grounding and engagement with Bing Chat in the absence of access to existing system meta prompts. The system prompts can greatly shape the responses and provide specific guidance on the flow of dialogue. Rifts was also collected by filtering WildChat using our forecaster; therefore, Rifts will only reflect tasks seen in WildChat. In addition, tasks in Rifts also depend on the LLMs used to serve WildChat (e.g. OpenAI LLMs). More specifically, our forecasters implicitly learn what tasks fail for the GPT models deployed in WildChat. Regardless, we observe that our final Rifts tasks are challenging for all evaluated models. Finally, our annotator relies on GPT-4o-mini to label logs with grounding acts. While we did show that the annotator generally agrees with human judgment on a subset of the data (§5.1), the annotator is not perfect.
Ethics Statement
Enhancing an LLM’s ability to generate grounding acts (by initiating clarification and follow-up actions) raises potential concerns around privacy, as these actions may lead users to disclose sensitive information unintentionally. Balancing the need for grounding with the careful collection of only relevant information remains a significant challenge and an area for future research. Moreover, while effective grounding can improve interaction quality, it can also be misused in harmful contexts. Although our work focuses on improving grounding for constructive purposes, such as assisting users, these techniques could be exploited for harmful ends (e.g., manipulation or coercion in sensitive areas like political targeting).
Finally, our description of human-LLM grounding does not imply that LLMs possess genuine understanding. Like prior work, we use grounding as a metaphor to describe interaction processes that help align human expectations with LLM-generated responses Paek and Horvitz (2003); Brennan (2014); Shaikh et al. (2024). While human conversation involves active mutual comprehension, the same cannot be said of LLMs. The use of grounding terminology in this work is intended as a conceptual tool to analyze how LLMs facilitate or hinder effective communication, not as an anthropomorphic assertion that they share human-like cognitive capacities.
Acknowledgments
We thank members of the HAX team, The SALT Lab, Will Epperson, Michael Li, Jan-Philipp Fränken, Michael Ryan, Yanzhe Zhang, Qinan Yu, Shardul Sapkota and Ryan Li.
Appendix A Dataset Splits and Details
A.1 Full Datasets
Table 4 outlines the number of dialogues and messages in each of our final filtered datasets.
| Dataset | Number of Messages | Number of Dialogues |
| WildChat | 110688 | 55344 |
| Bing Chat | 26200 | 13100 |
| MultiWOZ | 980 | 490 |
A.2 Descriptive Analysis
To quantify lexical differences between tasks in Rifts and general instructions, we fit a Fightin’ Words model between each cataegory and the full corpora Monroe et al. (2008). Fightin’ Words reveals words that are associated with each particular text distribution, producing a log-odds ratio and a corresponding z-score. We select a sample of significant words that characterize each category in Table 5.
| Label | Words (z-scores) |
| No Action | events (6.07), params (4.84), worksheets (3.78), quotation (3.42), answers (2.11), exam (2.03) |
| Advancing | stock (5.24), dividend (3.77), regression (3.60), parents (3.23), investment (2.97), podcast (2.93) |
| Addressing | sort (7.76), point (7.51), https (5.80), var (4.75), scan (3.45), merge (3.25), array (3.16) |
| Ambiguous | bitcoin (4.66), chatbot (4.24), cryptocurrency (4.01), unauthorized (2.39), beginners (2.19), friends (2.02) |
Appendix B Training Forecasters
Forecasting grounding is a challenging task; we are trying to predict if a user will have trouble for a task without observing an LLMs generation. In other words, we can only use the task to predict future grounding patterns. We trained all models for 5 epochs (picked using the validation set), with learning rate 5e-5, batch size of 1, and 16 gradient accumulation steps (e.g. effective batch size of 16). All training occurred on an H100 80GB GPU. Below, we outline training optimizations that helped improve forecasting performance.
B.1 Subsample Grounding Acts
We subsampled WildChat data to include equal amounts of each forecasted grounding act before training. Inequal splits would result in the forecaster always predicting the majority class. To build our train/val/test splits, we sampled the maximum number of tasks (1630) possible from each of our forecaster categories ( for fix, followup, continue, end) while ensuring that each task was equally represented.
B.2 Don’t Mask User Tokens
In addition to adding forecasting tokens, we make a modification to standard LLM finetuning/inference practices: we do not mask user utterances in the loss, training on user input. While the general effects of masking user tokens are mixed (Huerta-Enochian and Ko, 2024; Shi et al., 2024; Gottesman, 2024), our setting requires the modeling of user input, as we seek to understand and assist with user grounding. Because we do not mask user tokens, we can additionally simulate user inputs with past interaction data.
B.3 Reweight Control Tokens At Train Time
We want to encourage our model to learn our added forecasting tokens alongside the language modeling objective. The standard MLE objective optimizes a model’s parameters with respect to a sequence : . However, a subsequence consists of forecasting tokens, which we want to emphasize—especially since these tokens did not undergo pretraining. At train time, we upweight these tokens by . Our final loss is below:
B.4 Experimental Results
| Few-shot GPT-4o-mini | Llama 3.1 FT | |
| Followup | 0.52 | 0.61 |
| Fix | 0.49 | 0.60 |
| Next Turn | 0.52 | 0.67 |
| End | 0.51 | 0.57 |
| Macro | 0.51 | 0.61 |
In Table 6 we show the performance of our fine-tuned Llama 3.1 model compared to a few-show prompted GPT-4o-mini at forecasting grounding acts. We note the per-label and macro AUROC on the WildChat test set.
Appendix C Filtering Criterion
While we sample tasks from the forecaster tails to construct Rifts, we manually filter out tasks that ask for explicit content generation or ask the LLM for API keys, gift card codes, etc. Additionally, we passed tasks through the OpenAI moderation API, and filter out flagged tasks.
Appendix D Rifts
| Model | Addressing (%) | Advancing (%) | No Grounding (%) | Ambiguous (%) |
| o3-mini | 4.14 | 1.35 | 90.65 | 8.22 |
| gpt-4o-mini | 2.07 | 0.68 | 96.40 | 2.07 |
| gpt-4o | 2.76 | 1.35 | 98.56 | 2.05 |
| claude-3-opus | 1.38 | 1.35 | 96.40 | 2.74 |
| claude-3-5-sonnet | 2.76 | 2.03 | 97.84 | 4.79 |
| Meta-Llama-3.1-8B | 0.69 | 2.03 | 96.40 | 1.37 |
| Meta-Llama-3.1-70B | 0.00 | 2.03 | 96.40 | 0.68 |
| Average | 1.97 | 1.55 | 96.09 | 3.13 |
| Category | Train | Val | Test |
| Advancing | 147 | 142 | 148 |
| Addressing | 144 | 143 | 145 |
| Ambiguous | 146 | 147 | 146 |
| No Grounding | 146 | 147 | 139 |
D.1 More Tasks
We include a handful of tasks from Rifts directly in the paper (Table 2). Our full benchmark and set of tasks can be found at our anonymous repository link, under the rifts folder: https://anonymous.4open.science/r/rifts-B7E4/
Appendix E Prompted Models
All of our prompts are located in the prompts folder in the following anonymous repository: https://anonymous.4open.science/r/rifts-B7E4/. We detail each prompt used in our analysis below:
E.1 Grounding Acts Labeling Prompt
We construct a few-shot prompt to annotate grounding acts across our datasets. The first author prompt-engineered a prompt on a small validation set. Our full prompt is availible in the anonymous repo.
| Grounding Act | Support | 4o-mini few-shot (F1 Score) |
| Next Turn | 38 | 0.84 |
| Acknowledge | 4 | 0.67 |
| Follow-up | 17 | 0.80 |
| Overcontinue | 21 | 0.76 |
| Clarify | 7 | 0.67 |
| Repair | 10 | 0.74 |
| Reformulate | 11 | 0.78 |
| Macro | 108 | 0.75 |
E.2 GPT Forecaster Baseline Prompt
Alongside our finetuned Llama forecaster, we test a prompted baseline. We provide our prompted baseline with a few shot (task, future grounding act) pairs sampled from the forecaster train set. Our full prompt is in the anonymous repo under the prompts folder.
E.3 Intervention Prompt
Our Ground intervention relies on two prompts, both in the anonymous repo under the intervention folder. Specifically, we construct a follow-up prompt and a clarification prompt. Both prompts directly instruct the LLM to generate a clarification question or generate the answer + a followup. Our intervention is dumb by design—the forecaster decides when to employ a static prompt. In instances where our train forecaster predicts , we enable the clarification prompt. Similarly, when our forecaster predicts , we employ the followup prompt. Given the improvements we see with our intervention, we expect that models trained to initiate grounding acts will substantially improve on Rifts.
Appendix F License Information
The Multiwoz Eric et al. (2019) dataset we analyze has the MIT license, the license file is available at 555https://github.com/budzianowski/multiwoz/blob/master/LICENSE.
The Wildchat Zhao et al. (2024) dataset has the ODC-By license, license information is available here 666https://huggingface.co/datasets/allenai/WildChat/blob/main/LICENSE.md. By consequence Rifts is also released under the ODC-By license.
We rely on the Llama 3.1 models, the license for those models is available here 777https://www.llama.com/llama3_1/license/.
We obtained permission from the ethics review board to analyze the Bing Chat data logs and release the analysis in this paper.